OpenStudy (anonymous):

5.2x + 3.2 = 9.8 solve for x

OpenStudy (anonymous):

5.2x=6.6 (You subtract 3.2 to the other side) Divide. x=6.6/5.2 Answer x=1.269230769 or 1.27